I'm having a hard time understanding the solving of the equation?

Y=4-x
Y=x-4
Getting tired

Answer by mananth(16946)

Y=4-x
Y=x-4
since both are equal to Y
4-x=x-4
-2x=-8
/-2
x=4
Plug x=0 in any equation above
y=0